OpenStudy (anonymous):

mr. butterworth baked a cake in the shape of a cylinder. the cake had a diameter of 9 in. and a height of 5 in. he spread chocolate icing over the entire cake(except the bottom) how many square inches of icing did he use?

OpenStudy (andrewkaiser333):

chris from yahoo answers said this and i agree kinda The area will be: TOP OF CAKE =pi*r*r= 22/7*9/2*9/2=891/14=445.5/7=63.2 inches^2 SIDE OF CAKE =2*pi*d*h=2*22/7*9*5=180*11/7=1980/7= 282.8 inches^2 ADD TOGETHER =2425.5/7= 346.5 inches^2 May be slightly off because of mental arithmetic

OpenStudy (andrewkaiser333):

Did that help

OpenStudy (anonymous):

kinda. idk. its not in one of my options